OpenStudy (anonymous):

During a workout at the gym, you set the treadmill at a pase of 55.0 m/min. How many minutes will you walk if you cover a distance of 7500 ft?

OpenStudy (anonymous):

7500/(# of mile in feet) = x/55?

OpenStudy (anonymous):

7500/5280 = x/55, I think

OpenStudy (anonymous):

I have the answer; but I don't know how to show my work :-(

OpenStudy (anonymous):

5280/55 is your rate of feet per minute. Now divide 7500 by the rate at which you ran. That should give you 78.125, unless I'm mistaken.

OpenStudy (anonymous):

the answer is 42 mins

OpenStudy (anonymous):

How could it be less, having run more distance at the same rate?

OpenStudy (anonymous):

I don't understand the equation I got the same answer you have; but according to the book it's 42 mins

OpenStudy (anonymous):

oh gosh... is that 55.0 meters? I must have read it wrong, sorry. You have to do a meters to inch conversion and determine from there.

OpenStudy (anonymous):

convert the distance in feet to distance in meters

OpenStudy (anonymous):

x = how long did it take 7500ft * 0.3048m = 2286 meters rate * time = distance 55 * x = 2286

OpenStudy (anonymous):

thanks